OpenStudy (anonymous):

Of 100 students surveyed, 33 like orange soda, 55 like cola and 14 like both. How many did not like orange soda? 

OpenStudy (anonymous):

what do you think

OpenStudy (anonymous):

I am thinking it is either 67 or 69.

OpenStudy (anonymous):

55 like cola and 14 like both so i would go with 55

OpenStudy (anonymous):

I am not sure. Because when I solve it after making a venn diagram, I get 67. When I asked a tutor on a different site, he told me 69. I am not sure.

OpenStudy (amistre64):

lets think about this spose we have 2 groups, that have something in common A = 1+2 B = 2+3 yet we know that A and B make up the total amount given that the total amount is 6, what went wrong with adding A to B?

OpenStudy (amistre64):

Of 100 students surveyed, 33 like orange soda, 55 like cola and 14 like both. How many did not like orange soda? like orange = 33, or rather 19+14 like soda = 55, or rather 41+14 iof those that like orange, 33, only 14 of them like both, so 19 is left over that just like orange alone

OpenStudy (amistre64):

i hate these reverse negative questions .... not like orange is therefore 55 - 14 = 41 that is

OpenStudy (amistre64):

 and i assumed the total of the sets was 100 which is an error on my part as well 100 = 33 + 55 - 14 + (dont like either) 114 - 88 = (dont like either) 26 = (dont like either) so we have to count them as well as the 41 already accounted for 26 + 41

OpenStudy (amistre64):